OpenStudy (anonymous):

Can someone please help the first one is find the sum of the first 5 terms of the geometric series 7,21,63... the choices are A.91 B.280 C.847 D.2548 the second is vernas starting salary is 32,000 a year she will get a 4% raise annually how much money will she have made after 7 years? A.173,322.32 B212,255.21 C.252,745.42 D.294,855.24 and the last one which of the following is not a geometric sequence? A.1,2,4,8 B. 3,1,-1,-3 C. 9,14,19,24 D. 8,13,18,23 i did the rest of the questions but i got stuck on these three my teacher was absent and I had a sub so we didnt learn this.

OpenStudy (anonymous):

@Lena772

OpenStudy (lena772):

Ok well for the first question each number in the sequence is being multiplied by 3. They only give you the first 3 terms. 7, 21, 63. 63*3=189 189*3= 567 So those are the next two terms. 7+21+63+189+567=?

OpenStudy (anonymous):

847?

OpenStudy (lena772):

Yup, C. :)

OpenStudy (lena772):

Money is being raised 4% each year. So 104% of the current salary is the new salary 104/100*32000=33280 (salary after first year) 104/100*33280= 34,611.20 (salary after second year) Can you continue from here to find her salary after the seventh year?

OpenStudy (anonymous):

yes i understand what about the last question

OpenStudy (lena772):

A geometric sequence is one that goes from term to term by the subtraction or addition of the same value.

OpenStudy (lena772):

So A would not be a geometric sequence. :)

OpenStudy (anonymous):

oh ok i understand thanks so much oh hold on sorry one last question before you go if you dont mind

OpenStudy (lena772):

Sure?

OpenStudy (anonymous):

find the s25 for 3+7+11+15+... the choices are A.1,263 B.1,267 C.1,271 D. 1,275

OpenStudy (lena772):

s25?

OpenStudy (lena772):

you mean the 25th number in the sequence?

OpenStudy (anonymous):

yea thats what the question asked i think so it doesnt clarify

OpenStudy (lena772):

So each time where adding on a number that is increased by a certain value, can you see which value this is?

OpenStudy (anonymous):

4?

OpenStudy (lena772):

YES!

OpenStudy (lena772):

Good

OpenStudy (lena772):

So we are finding the SUM of the first 25 numbers of the sequence.

OpenStudy (lena772):

3+7+11+15+19+23+27+31+35+39+43+47+51+55+59+63+67+71+75+79+83+87+91+95+99

OpenStudy (lena772):

Count the numbers, make sure there's 25

OpenStudy (lena772):

3+7+11+15+19+23+27+31+35+39+43+47+51+55+59+63+67+71+75+79+83+87+91+95+99=?

OpenStudy (anonymous):

1257?

OpenStudy (anonymous):

hey before we countinue for the one with the salary i keep get more than my anwser choices idk why/?

OpenStudy (lena772):

No, not correct.

OpenStudy (lena772):

I will help you with that after we get this one done.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its 1275

OpenStudy (anonymous):

sorry i flipped the last number

OpenStudy (lena772):

yup, D!

OpenStudy (anonymous):

okay and about the third one i keep getting 42107?

OpenStudy (lena772):

104/100*32000=33280 (salary after first year) 104/100*33280= 34,611.20 (salary after second year) 104/100*34611.2=35995.65 (salary after third year) 104/100*35995.65=37435.48 (salary after fourth year) 104/100*37435.48=38932.90 (salary after fifth year) 104/100*38932.90=40490.22 (salary after sixth year) 104/100*40490.22=42109.83 (salary after seventh year) 32000+33280+34,611.20+35995.65+37435.48+38932.90 +40490.22+42109.83=?

OpenStudy (anonymous):

oh wow okay thanks so i was doing it right i just read the number wrong thanks so much for all your help ill make sure to fan you ...

OpenStudy (lena772):

You did it right, you just didn't add her salaries together to see how much she made! So what is the sum of all of her salaries?

OpenStudy (anonymous):

i got 294855.28?

OpenStudy (anonymous):

i mean .24

OpenStudy (anonymous):

i have to go but thank you so much for your help .

OpenStudy (lena772):

You're welcome, goodbye.

OpenStudy (lena772):

Ranga is right

OpenStudy (ranga):

@Lena772 and @Karina1: You can get the answers quickly if you use the standard formula for the sum of the first n terms of a geometric series and an arithmetic series:\[\text {Geometric Series:}~~a + ar + ar^2 + ar^3 + ... + ar^{n-1}\]\[\quad \quad \text {Sum = } ~ a \frac{ 1 - r^n }{ 1 - r }\]\[\text {Arithmetic Series}:~~a _{1} + (a _{1} + d) + (a _{1} + 2d) +...+ (a _{1} + (n-1)d)\]\[\quad \quad \text {Sum} = \frac{ n }{ 2 }(a _{1} + a _{n}). \quad \text { Last term } a_{n} = a_{1} + (n - 1)d\] Verna's starting salary is 32,000 a year she will get a 4% raise annually how much money will she have made after 7 years? The teacher's salary starts at 32000 and gets multiplied by 1.04 each year (4% raise). It is a geometric series. What is the sum after 7 years? a = 32000; r = 1.04; n = 7; Sum = a * (1 - r^n) / (1 - r) = 32000(1 - 1.04^7) / (1 - 1.04) = 252,745.42 For the problem: find the S25 for 3+7+11+15+... the choices are A.1,263 B.1,267 C.1,271 D. 1,275 It is an arithmetic progression with the first term a1 = 3, common difference d = 4 and total number of terms n = 25. The nth term an is given by the formula: an = a1 + (n-1)d an = 3 + (25-1)(4) = 3 + 24*4 = 99 Sum = n/2 * (a1 + an) = 25/2 * (3 + 99) = 1275. Choice D. Which of the following is not a geometric sequence? A.1,2,4,8 B. 3,1,-1,-3 C. 9,14,19,24 D. 8,13,18,23 A is geometric sequence because each term gets MULTIPLIED by the same number, which in this case is 2. B is arithmetic sequence because each term DIFFERS from the previous by the same amount and in this case it is -2. C is arithmetic sequence because each term DIFFERS from the previous by the same amount and in this case it is 5. D is arithmetic sequence because each term DIFFERS from the previous by the same amount and in this case it is 5. B,C,D are NOT geometric sequences. A is a geometric sequence.